THEORETICAL STUDY ON FERMI RESONANCE OF THE EXCESS PROTON VIBRATION IN BINARY CLUSTERS

\underline{JER-LAI KUO}, JAKE TAN, Institute of Atomic and Molecular Sciences Academia Sinica, Taipei 10617, Taiwan; RYUNOSUKE SHISHIDO, ASUKA FUJII, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

In protonated binary clusters, location and magnitude of delocalization of the excess proton is shown to correlate to the difference of the proton affinities of the two components^{a}. Theoretical calculations on a series of protonated binary clusters have shown that Fermi resonance can occur in clusters containing amines. For examples, in (CH_{3})_{3}N-H^{+}-X, the N-H^{+} stretching frequency is high than the first overtone of N-H^{+} bending modes. As the N-H^{+} stretching frequency decreases by increasing the proton affinity of X, Fermi resonance inevitably shall occur. At least both N-H^{+} bending modes seem to couple with the N-H^{+} stretching mode; we performed multi-dimensional mode analyses to understand this coupling.

aJ. R. Roscioli, L. R. McCunn, M. A. Johnson Science316, 249 (2007).